OpenStudy (anonymous):

Write the slope-intercept form of the line which passes through (4, 2) with a slope of 1/2. A. y = 1/2 x + 2 B. y = 1/2 x C. y = 1/2 x + 4

OpenStudy (anonymous):

use point slope form

OpenStudy (nubeer):

m = 1/2 x1 = 4, y1 = 2 now use this formula. y-y1=m(x-x1) just plug in the values .. you will get the equation